# Targeted Nakamoto: A Bitcoin Protocol to Balance Network Security and Energy Consumption

### v2

@@ 5.1 The block reward adjustment allocations
One of the T constraints is that $Targeted$ should not change the order in which transactions are appended to blocks. Huberman et.al. [HLM21] demonstrate that Nakamoto induces an equilibrium ordering of transaction fees and transactions on the blockchain that reflects transactors’ relative time preferences. Transactions are appended to blocks in the order of the size fee offered and the ranking of fees match the ranking of transactors’ time-preference for executing their transaction. [^29] See Theorem 1 and Proposition 2. Note that the result relies on parametric assumptions. The reader can consult Huberman et.al. [HLM21] for details. It is a constrained efficient outcome in the sense that if agents A and B send their respective transactions to the mempool at the same time and agent A incurs a higher cost for delay in completing its transaction than agent B, agent A’s transaction will be appended to the blockchain ahead of agent B’s transaction. Targeted preserves the same ordering of transactions by adjusting the coinbase and each transaction fee by the same percentage. $Fee_{f,B_{n}}$ denotes the fee paid by the $f$th transaction appended to block $B_{n}$. Denoting $\xi=(p^{M}_{B_{n}}-p_{B_{n}})/p_{B_{n}}$, the ordering $Fee_{i,B_{n}}>Fee_{j,B_{n}}$ implies the ordering $\xi Fee_{i,B_{n}}>\xi Fee_{j,B_{n}}$, i.e. the ordering of transaction fees sent to the miner is unchanged. $R_{n}$ denotes the coinbase portion of the block reward. The miner’s block reward is $p^{M}_{B_{n}}=(1+\xi)\big{[}\sum_{f}Fee_{f,B_{n}}+R_{n}\big{]}$. When $\xi<1$ the quantity $\xi p_{B_{n}}$ UTXO’s is burned. When $\xi>1$ the quantity $\xi p_{B_{n}}$ UTXO’s are minted (above the scheduled coinbase).
@@ 5.3 Policy to control hashrate
The Hashrate Control Algorithm.
A ceiling or floor on the miner’s block reward is adjusted after the end-block $E_{q}$ of each epoch $q$ as a function of $D/T$. There are four cases to consider.
Case 1: The puzzle adjustment at block $E_{q}$ implies that $D/T$ crosses $\frac{D}{T}(N_{UB})$ from below.
@@
Case 2: The puzzle adjustment at block $E_{q}$ implies that $D/T$ crosses $D\frac{D}{T}(N_{UB})$ from above and is in the in the target hashrate interval.
For each subsequent epoch until $\text{ceil}_{\gamma}\leq\overline{P_{B}}(\gamma)$, the ceiling is increased by the formula $\text{ceil}_{q+1}\leftarrow(1-\tau)\text{ceil}_{q}$ Thereafter, the ceiling is removed.
Case 3: The puzzle adjustment at block $E_{q}$ implies that $D/T$ crosses $\frac{D}{T}(N_{LB})$ from above.
A floor on the miner’s block reward is imposed. For each block in epoch $q+1$, payment to miner is capped at $\text{floor}_{q}=(1-\tau)\overline{P_{B}}(q)$, where $\overline{P_{B}}(q)$ is the median block reward in Epoch $q$.
For each subsequent epoch until $\frac{D}{T}(N_{LB})$ is crossed from below, the floor is raised by the formula $\text{floor}_{q+1}\leftarrow(1-\tau)\text{ceil}_{q}$.
Case 4: The puzzle adjustment at block $E_{q}$ implies that $D/T$ crosses $\frac{D}{T}(N_{LB})$ from below and is in the in the target hashrate interval.
For each subsequent epoch until $\text{floor}_{\gamma}\geq\overline{P_{B}}(\gamma)$, the floor is reduced by the formula $\text{floor}_{q+1}\leftarrow\tau\text{floor}_{q}$ Thereafter, the floor is removed.  [^31] The choice of $\tau$ and $(1-\tau)$ as ceiling and floor adjustment parameters is arbitrary. The optimization of the adjustment parameters is left for future research.
@@ 6 Dynamic Adjustment of Hashrate
This section evaluates the dynamic response of hashrate to changes in the value of the miner’s block reward which occur when hashrate is outside the target interval. The dynamic adjustment of hashrate to a change in the miner’s block reward can be evaluated when hashrate adjusts to maintain a profit $\phi$ per target time $T^{*}$. Consider an epoch $q$ where the total block reward is constant at $p^{M}_{B}$ and the puzzle difficulty adjustment at the end of epoch $q-1$ achieves the target growth rate when hashrate is $N$, and $N$ is the equilibrium hashrate when the miner’s block reward is $p^{M}_{B}$. Suppose a ceiling or floor is imposed at the end of epoch $q-1$ resulting in a marginal change in the miner’s block reward equal to $\Delta p$ (i.e. holding fixed the exchange rate, $e$, and hashrate cost $c$). Let $\Delta_{T}$ denote the distance between the actual blockchain growth rate and the target blockchain growth rate induced by a change in hashrate in epoch $q$, denoted $\Delta N$.
@@ 7.1 The Targeted Monetary Policy
Figure 5 depicts the operation of the Targeted monetary policy at block $B_{n}$. The timing sequence is denoted by the circled numbers. The first step is the execution of the transactions. The adjustment to a transaction sender’s UTXO value is determined by the difference between its account value in the UTXO pool and its account value in the Remainder Accounts. The second step is the payment of the miner block reward. If $P^{M}_{B_{n}}<p^{M}_{B_{n}}$, UTXO value of $q$ is minted to cover the payment. The third step is to send of UTXO value $q$ to the UTXO Pool if $P^{M}_{B_{n}}<p^{M}_{B_{n}}$ or else burn UTXO value up to $q$ in the UTXO pool and increase the value of the Remainder Accounts by the residual. Adjustments to the address account values in the UTXO Pool and the Remainder Accounts are made.383838 Steps 2 and 3 can be simultaneous. When $q<0$ the minting of UTXO value is reduced by the amount of value in the UTXO pool, which is sent to the miner from the UTXO pool. This is implicit in the code in Appendix B.
@@ Appendix A Code for The Hashrate Control Algorithm
Listing 1: Python code for Hashrate Control Algorithm
@@
self.rewards = [] # To store the rewards of each block within an epoch
@@
return statistics.median(sorted_rewards)
@@
def __init__(self, tau):
@@
self.DT = None # Puzzle difficulty/Blockchain growth rate; get this value from blockchain source code
self.DT_N_UB = None # Upper boundary; this needs to be set from outside
self.DT_N_LB = None # Lower boundary; this also needs to be set from outside
@@
# Case 1
if self.DT > self.D_N_UB:
self.epochs.append(Epoch())
self.epochs[-1].ceil = self.tau * P_B_median
# Cases 2 and 4
elif self.DT < self.DT_N_UBand self.DT > self.DT_N_LB:
self.epochs.append(Epoch())
if last_epoch.ceil != None:
if last_epoch.ceil <= P_B_median:
self.epochs[-1].ceil = (1 - self.tau) * last_epoch.ceil
@@
self.epochs[-1].ceil = None
elif last_epoch.floor != None:
if last_epoch.floor >= P_B_median:
self.epochs[-1].floor = self.tau * last_epoch.floor
@@
self.epochs[-1].ceil = None
@@
self.epochs[1].floor = None
# Case 3
elif self.DT < self.D_N_LB:
self.epochs.append(Epoch())
self.epochs[-1].floor = (1 - self.tau) * P_B_median
@@ Appendix B Code for the Targeted Monetary Policy
Listing 2: Python code for Targeted monetary policy

### v3

@@ 5.1 The block reward adjustment allocations
One of the T constraints is that $Targeted$ should not change the order in which transactions are appended to blocks. Huberman et.al. [12] demonstrate that Nakamoto induces an equilibrium ordering of transaction fees and transactions on the blockchain that reflects transactors’ relative time preferences. Transactions are appended to blocks in descending order of the size fee offered and the ranking of fees match the ranking of transactors’ time-preference for executing their transaction. [^29] See Theorem 1 and Proposition 2. Note that the result relies on parametric assumptions. The reader can consult Huberman et.al. [12] for details. It is a constrained efficient outcome in the sense that if agents A and B send their respective transactions to the mempool at the same time and agent A incurs a higher cost for delay in completing its transaction than agent B, agent A’s transaction will be appended to the blockchain ahead of agent B’s transaction. Targeted preserves the same ordering of transactions by adjusting the coinbase and each transaction fee by the same percentage. $Fee_{f,B_{n}}$ denotes the fee paid by the $f$th transaction appended to block $B_{n}$. Denoting $\xi=(p^{M}_{B_{n}}-p_{B_{n}})/p_{B_{n}}$, the ordering $Fee_{i,B_{n}}>Fee_{j,B_{n}}$ implies the ordering $\xi Fee_{i,B_{n}}>\xi Fee_{j,B_{n}}$, i.e. the ordering of transaction fees sent to the miner is unchanged. $R_{n}$ denotes the coinbase portion of the block reward. The miner’s block reward is $p^{M}_{B_{n}}=(1+\xi)\big{[}\sum_{f}Fee_{f,B_{n}}+R_{n}\big{]}$. When $\xi<1$ the quantity $\xi p_{B_{n}}$ UTXO’s is burned. When $\xi>1$ the quantity $\xi p_{B_{n}}$ UTXO’s are minted (above the scheduled coinbase).
Crucially, this method of proportional adjustment of the fee paid to the miner does not alter the Nash equilibrium of the pattern of fees offered by transactors. This is readily apparent by considering the incentive for $A$ or $B$ to alter their fee after a proportional adjustment $\xi$ of the miner’s fee. Miners will still have an incentive to append transactions to blocks in ascending order of fees. If $B$ does not change its fee, $A$’s incentives are unaffected. It still must offer a higher fee than $B$ to secure a position in the queue of transactions ahead of $B$. Similarly, if $B$ was unwilling to offer a higher fee than $A$ before the adjustment, it not have an incentive to do so after the adjustment.
@@ 5.3 Policy to control hashrate
The Hashrate Control Algorithm A ceiling or floor on the miner’s block reward is adjusted after the end-block $E_{q}$ of each epoch $q$ as a function of $D/T$. There are four cases to consider.
Set parameters $\tau,\gamma\in[0,1]$.
Case 1: Hashrate is above the upper bound of the hashrate interval, $\frac{D}{T}(N_{UB})$.
@@
Case 2: The puzzle adjustment at block $E_{q}$ implies that $D/T$ is below the lower bound of the target hashrate interval, $\frac{D}{T}(N_{LB})$.
A floor on the miner’s block reward is imposed. For each block in epoch $q+1$, payment to miner is capped at $\text{floor}_{q}=(1+(1-\tau))\overline{P_{B}}(q)$, where $\overline{P_{B}}(q)$ is the median block reward in Epoch $q$.
For each subsequent epoch until $\frac{D}{T}(N_{LB})$ is crossed from below, the floor is raised by the formula $\text{floor}_{q+1}\leftarrow(1+(1-\tau))\text{ceil}_{q}$.
Case 3: The puzzle adjustment at block $E_{q}$ implies that $D/T$ is inside the target hashrate interval.
For each subsequent epoch until $\text{ceil}_{q}\leq\overline{P_{B}}(q)$, the ceiling is increased by the formula $\text{ceil}_{q+1}\leftarrow(1+(1-\gamma)\text{ceil}_{q}$ Thereafter, the ceiling is removed.
For each subsequent epoch until $\text{floor}_{q}\geq\overline{P_{B}}(q)$, the floor is reduced by the formula $\text{floor}_{q+1}\leftarrow\gamma\text{floor}_{q}$ Thereafter, the floor is removed.
 [^31] The formulas $\tau$ and $(1+(1-\tau))$ ensure symmetry in the adjustment of the ceiling and the floor, i.e. that the ceiling and floor adjust by the same percentage each epoch.
@@ 6 Dynamic Adjustment of Hashrate
This section evaluates the dynamic response of hashrate to changes in the value of the miner’s block reward which occur when hashrate is outside the target interval. The dynamic adjustment of hashrate to a change in the miner’s block reward can be evaluated when hashrate adjusts to maintain a profit $\phi$ per target time $T^{*}$. [^32] The assumption of a constant profit rules out the possibility of a dynamic strategy to influence hashrate by a miner coalition. Arguably, $\phi$ internalizes the exercise of market power.
Consider an epoch $q$ where the total block reward is constant at $p^{M}_{B}$ and the puzzle difficulty adjustment at the end of epoch $q-1$ achieves the target growth rate when hashrate is $N$, and $N$ is the equilibrium hashrate when the miner’s block reward is $p^{M}_{B}$. Suppose a ceiling or floor is imposed at the end of epoch $q-1$ resulting in a marginal change in the miner’s block reward equal to $\Delta p$ (i.e. holding fixed the exchange rate, $e$, and hashrate cost $c$). Let $\Delta_{T}$ denote the distance between the actual blockchain growth rate and the target blockchain growth rate induced by a change in hashrate in epoch $q$, denoted $\Delta N$.
@@ 7.1 The Targeted Monetary Policy
Figure 5 depicts the operation of the Targeted monetary policy at block $B_{n}$. The timing sequence is denoted by the circled numbers. The first step is the execution of the transactions. The adjustment to a transaction sender’s UTXO value is determined by the difference between its account value in the UTXO pool and its account value in the Remainder Accounts. The second step is the payment of the miner block reward. If $P^{M}_{B_{n}}<p^{M}_{B_{n}}$, UTXO value of $q$ is minted to cover the payment. The third step is to send of UTXO value $q$ to the UTXO Pool if $P^{M}_{B_{n}}<p^{M}_{B_{n}}$ or else burn UTXO value up to $q$ in the UTXO pool and increase the value of the Remainder Accounts by the residual. Adjustments to the address account values in the UTXO Pool and the Remainder Accounts are made.393939 Steps 2 and 3 can be simultaneous. When $q<0$ the minting of UTXO value is reduced by the amount of value in the UTXO pool, which is sent to the miner from the UTXO pool. This is implicit in the code in Appendix B.
@@ Appendix A Code for The Hashrate Control Algorithm
 1: Python code for Hashrate Control Algorithm
@@
self.rewards = [] # To store the rewards of each block
within an epoch
self.hashrates = [] # To store the hashrates of each block
within an epoch
@@
self.difficulty = None
self.elapsed_time = None # T_n
@@
return np.median(sorted_rewards)
def average_hashrate(self):
sorted_hashrates = sorted(self.hashrates)
return np.mean(sorted_hashrates)
@@
def __init__(self, tau, gamma, upper_bound, lower_bound, initial_difficulty):
@@
self.DT = None # Puzzle difficulty/Blockchain growth rate;
get this value from blockchain source code
self.DT_N_UB = upper_bound # Upper boundary
self.DT_N_LB = lower_bound # Lower boundary
@@
def adjust_hashrate(self, hashrate, epoch_idx, e_current, P_current, efficiency_current, electricity_cost_current):
alpha1 = None
alpha2 = None
alpha3 = None
alpha4 = None
# Found through regression on past values
epoch = self.epochs[epoch_idx]
log_eP_current = np.log(1+ (e_current * P_current))
log_eff_current = np.log(1+efficiency_current)
log_electricity_cost_current = np.log(1+electricity_cost_current)
# Put it all together:
[…]
predicted_log_hashrate = (
alpha1
+ alpha2 * log_eP_current
+ alpha3 * log_eff_current
+ alpha4 * log_electricity_cost_current
[…]
new_hashrate = np.exp(predicted_log_hashrate)
epoch.add_hashrate(new_hashrate)
return new_hashrate
@@
N_n = last_epoch.average_hashrate()
new_epoch = Epoch()
BLOCKS_PER_EPOCH = 14
T_STAR = BLOCKS_PER_EPOCH * 10 * 60 # 1,209,600 for 2016 blocks
D_n = last_epoch.difficulty
T_n = (BLOCKS_PER_EPOCH * D_n * 2**32) / N_n
last_epoch.elapsed_time = T_n
T_hat = max(T_STAR /4, min(T_n, 4 *T_STAR))
D_next = D_n * (T_STAR / T_hat)
new_epoch.difficulty = D_next
# Case 1: Hashrate too high - need to decrease rewards
if self.DT > self.DT_N_UB:
#("Case 1: Hashrate above upper bound")
new_epoch.ceil = (self.tau) * P_B_median
new_epoch.floor = None # Clear any existing floor
# Case 2: Hashrate within bounds - gradual adjustment
elif self.DT_N_LB < self.DT < self.DT_N_UB:
# Previous epoch had a ceiling
if last_epoch.ceil is not None:
if last_epoch.ceil >= self.DT:
# Gradually relax the ceiling
new_epoch.ceil = (1+self.gamma) * last_epoch.ceil
@@
# Remove ceiling if it’s no longer needed
new_epoch.ceil = None
# Previous epoch had a floor
elif last_epoch.floor is not None:
if last_epoch.floor <= self.DT:
# Gradually relax the floor
new_epoch.ceil = (self.gamma) * last_epoch.floor
@@
# Remove floor if it’s no longer needed
new_epoch.floor = None
@@
new_epoch.floor = last_epoch.floor
new_epoch.ceil = last_epoch.ceil
# Case 3: Hashrate too low - need to increase rewards
elif self.DT < self.DT_N_LB:
#("Case 3: Hashrate below lower bound")
new_epoch.floor = (1+(1-self.tau)) * P_B_median
new_epoch.ceil = None # Clear any existing ceiling
self.epochs.append(new_epoch)
@@ Appendix B Code for the Targeted Monetary Policy
 2: Python code for Targeted monetary policy